# Directional Diffusion-Style Code Editing Pre-training

### v1

@@ Abstract
Code pre-trained models have shown promising effectiveness in various software engineering tasks. Among these tasks, many tasks are related to software evolution and/or code editing. However, existing code pre-trained models often overlook the real-world code editing data and the evolutionary nature of the editing process. In this paper, to simulate the step-by-step code editing process of human developers, we propose DivoT5, a pre-trained model based on directional diffusion at the data level. In DivoT5, we adopt two categories of pre-training tasks. The first category is mask and denoising tasks augmented with a diffusion direction representing code evolution. That is, we first apply a noising process to the code snippets before evolution, and then ask the pre-training process to restore the snippets with noise into the code snippets after evolution. The second category is tasks aiming to reinforce the evolutionary direction. That is, we first generate various intermediate versions for each pair of snippets before and after evolution, and then ask the pre-training process to transform the intermediate versions into the snippet after evolution for each pair. We evaluate DivoT5 for two code-editing scenarios (including a number of tasks) and one non-editing scenario using five downstream tasks. Given each downstream task, we fine-tune the pre-trained DivoT5 on the dataset for the downstream task and evaluate the effectiveness of DivoT5. Our experimental results show that DivoT5 achieves state-of-the-art (SOTA) performance on most tasks in comparison to models of the same scale (220M), large scale (770M) models in fine-tuning, and billion-scale (6.7B, 8B, ChatGPT) models in few-shot settings. For one code-editing task (i.e., automated code review), DivoT5 pre-trained on top of CodeT5-small (60M) can even outperform CodeT5-base (220M) and other pre-trained models with 220M parameters except for DivoT5 pre-trained on top of CodeT5-base (220M).
Index Terms:
Pre-training, code editing, software evolution
@@ I Introduction
The code of software systems often requires continuous maintenance and updates to enhance functionality, adapt to new requirements, and fix vulnerabilities [1, 2, 3, 4, 5], which are collectively termed “code evolution”. Throughout the code evolution process, many tasks require generating new code by editing existing old code, such as code refinement [6, 7] and automated code review [3, 8]. These tasks related to code editing constitute an essential part of the code evolution process. To address tasks related to code editing, many studies have utilized pre-trained code models, such as CodeT5 [9], given their proven ability to effectively encode semantic information of code. However, directly applying these pre-trained models to code editing tasks often fails to achieve satisfactory performance because they do not adequately consider the contextual information of code evolution. Contextual information offers guidance and reference for code evolution, which is crucial for proper modification of the code.
@@
Recently, several studies [8, 10, 7, 11] propose pre-training tasks specifically designed for code editing tasks and learning contextual information of code evolution by fine-tuning pre-trained models. These models consider code editing through two ways: masked language modeling (MLM) with code changes [7, 11] and denoising auto-encoding (DAE) with controlled noises [10, 8]. The first way represents code changes as the output sequence of the diff tool and applies the MLM pre-training tasks to understand the semantics of code changes. This way has shown promising results in understanding code changes. However, in real development scenarios, code evolution typically involves generating new code based on existing old code, rather than understanding the code change. Therefore, using code changes as input to generate corresponding descriptions (e.g., commit messages [11]) or classification labels (e.g., whether the code improves [7]) does not align well with such development scenarios for code generation. Moreover, the granularity they use is at the code line level [12] as output by the diff tool, also posing a challenge in considering fine-grained changes (i.e., token-level) during the generation process  (i.e., the first challenge lies in reflecting fine-grained edits akin to those made by human developers.). In the second way, noise is introduced to the old code based on the probability distribution derived from real code changes in downstream datasets, and the model employs the DAE pre-training tasks to restore the old code from the noisy input [8]. However, the noise-disrupted code merely reflects the overall probability of code being edited, without considering the varying contexts of the different code. Consequently, there exists a substantial gap between the disrupted code and the real-world imperfect code written by humans  (i.e., the second challenge lies in leveraging the context of code edits made by human developers). In the development scenarios, code evolution typically involves a series of modifications to one file, making it a mutli-step editing process rather than a one-off effort. Instead of focusing on only what the code changes are or what the actual change probabilities might be, how the code is gradually edited to achieve the final evolutionary goal should be also be considered. However, existing code editing pre-trained models only utilize the before-and-after code pairs directly, neglecting the gradual evolution of code in real-world scenarios and its characteristic of incremental editing  (i.e., the third challenge lies in capturing the incremental nature of code editing performed by human developers.).
@@
In this paper, we introduce DivoT5, a diffusion-style pre-trained model. To tackle the above challenges, the key idea of DivoT5 is to simulate the incremental nature of human code editing in the pre-training step. It draws inspiration from diffusion models, which excel at handling incremental transformations in continuous data domains, generating high-quality and diverse images or audio through a gradual multi-step denoising process [13, 14, 15]. Similarly, code editing involves making a series of iterative changes to fulfill the final requirements, making it naturally aligned with the step-by-step refinement principles of diffusion models. However, DivoT5 does not directly implement diffusion processes. It utilizes its data generation principles and focuses on simulating the gradual code editing steps that developers typically follow, allowing it to produce realistic and contextually relevant code modifications. To distinguish it from traditional diffusion methods, we refer to this approach as “directional diffusion.”
@@
Figure 1 illustrates the schematic of directional diffusion. We set the starting point (i.e., $X_{T}$) of the directional diffusion strategy as the old code, design three mask strategies as introducing artificial noise (i.e., $X_{T+1}$) to understand the evolutionary direction, and treat the intermediate states (i.e., $X_{t}$) of the editing process as evolutionary noise that reinforces the direction of evolution. Specifically, $X_{T+1}$ to $X_{T}$ represents the process of removing artificial noise (e.g., replacing $p_{1}$ and $v_{1}$ to $|unk|$ in Figure 1), while $X_{t}$ to $X_{t-1}$ represents the process of removing evolutionary noise (e.g., modifying the $f_{11}$ in the Figure 1). To fit the fine-grained editing scenarios, DivoT5’s pre-training takes the old code as input and incorporates token-level diff information when generating the complete new code. Rather than relying on line-level diff information to understand code changes, DivoT5 leverages token-level diff information to help the model discern which parts of the code need modification and which parts should remain unchanged (i.e., a mask strategy applied from $X_{T}$ to $X_{T+1}$ in Figure 1). To leverage and enrich the context of the real-world code edits, we design two additional mask strategies based on real code change histories. Learning from real code changes means that the model needs to automatically update old versions of code (which may contain bugs or require new features) to new versions that meet new requirements. These two masking strategies (applied from $X_{T}$ to $X_{T+1}$ in Figure 1) simulate potential issues with code quality or completeness in real code context by introducing varying degrees of random artificial noise into the old code. This requires the model to understand the context of the specific code and make modifications based on the provided guidance, rather than merely restoring the original version from the corrupted code. To learn the characteristics of incremental code evolution, we design a novel evolutionary direction reinforcement strategy by utilizing intermediate data in the code editing process. Intermediate versions (i.e., $X_{t}$ ($0<t\leq T$) in Figure 1) of the code that is not fully modified are seen as still containing evolutionary noise. We utilize intermediate data from the incremental evolution process to train the DivoT5 model, providing clear directional and more detailed guidance information to reinforce the evolutionary direction. Due to the limitations of applying existing diffusion models to discrete textual data, which typically use multi-step denoising at continued embedding levels and whose performance is constrained by the mapping between continuous embeddings and discrete texts, we apply the above strategies at the data level (i.e., conducting the diffusion process on data rather than embeddings). This model ensures the fluency and reliability of the generated results by directing each version towards the final version ($X_{0}$), even with the addition of artificial noise (i.e., constructing training samples $<X_{t}$, $X_{0}>$ and generating $X_{0}$ from any given $X_{t}$, where $10<t\leq T+$ 1). In other words, we adopt the encoder-decoder training framework [9] while incorporating the diffusion concept to emphasize the evolutionary direction of code editing. Rather than based on the full diffusion model structure, we draw on its style to enhance step-by-step refinement, ensuring the same computational efficiency as the encoder-decoder model [9, 8, 7] during training and inference.
@@
To evaluate DivoT5, we fine-tune DivoT5 on three scenarios (i.e., NL-guided code editing scenario, code-only editing scenario, and non-editing scenario) using five benchmark datasets. The evaluation results demonstrate that DivoT5 outperforms the existing state-of-the-art (SOTA) code pre-trained models for all five datasets. For example, DivoT5-base (220M) achieves a significant improvement in exact match performance compared to the CodeT5-base (220M) model, exceeding a 28.87% increase on the automated code review dataset (from 34.46% to 44.41%). Furthermore, DivoT5-base (220M) significantly outperforms the fine-tuned CodeT5-large (770M) model and billion-scale models like DeepSeekCoder-6.7B-instruct, Meta-Llama-3.1-8B-instruct, and ChatGPT-gpt-3.5-turbo in few-shot inference. The small version of DivoT5 (60M) also achieves a good performance in this task, surpassing CodeT5-base (220M) by 5 percentage points (from 34.46% to 40.16%) in the exact match metric. In addition, we evaluate the effectiveness of each pre-training task. The ablation results demonstrate the contributions of the designed pre-training tasks to the final performance.
@@ II-A Code-Editing Tasks and Code-Editing Related Tasks
During the daily development process, code editing is necessary to maintain the overall functionalities of a software project. In this paper, we differentiate two categories of tasks that are associated with code editing. The first category is tasks involving an editing process to change a code snippet, and we refer to this category as code-editing tasks [6, 3, 10, 8]. The second category is tasks that do not directly edit code but may use code-editing information for other purpose, and we refer to this category as code-editing related tasks [11, 7]. Obviously, it is helpful to automate both categories of tasks during the code maintenance process.
Based on whether natural language guidance is required during the code-editing process, we can further divide code-editing tasks into two types. We denote the first type as the NL-guided code editing scenario, such as automated code review [3, 8], NL-based code refinement [7]. Developers provide feedback and improvement suggestions by inspecting and evaluating the code, and the natural language represents the direction for code enhancement. The new code should align with this direction. We denote the second type as the code-only editing scenario, such as code refinement [16]. For the code refinement task, the new code should address the defects present in the old code. In this paper, we primarily consider these two types of code-editing tasks.
@@ II-B Code Pre-trained Models
When applying existing pre-training techniques [17, 18, 19, 20] to code data, there are two common pre-training tasks, which are referred to as masked language modeling (MLM) [21, 9] and denoising auto-encoding (DAE) [22]. The basic MLM technique refers to randomly masking a certain proportion of tokens in the input and then predicting the masked tokens. CodeBERT [21] and GraphCodeBERT [23] apply the MLM technique to its code data to enhance the model’s understanding of code. CodeT5 [9] employs the MLM technique to sequence-to-sequence generation tasks (referred to as SMLM) to enhance the model’s ability to generate code. Additionally, to recognize identifiers within the code, CodeT5 also introduces a variant of SMLM called IMLM (Identifier Masked Language Modeling). In IMLM, only the identifiers in the source code are masked, while the rest of the tokens are left unchanged, allowing the model to generate code while specifically paying attention to the discernment of identifiers. DAE is a self-supervised learning approach that enhances model generalization by introducing noise to the input and training the model to reconstruct the original input. In PLBART [22], DAE is introduced during the pre-training stage to train the model in code understanding and generation capabilities. Specifically, PLBART incorporates three noise functions: token masking, token deletion, and token infilling. These functions are applied with a certain probability to disrupt the old code.
@@ II-C Code Editing Pre-trained Models
Code editing pre-trained models are typically built upon existing code pre-trained models, taking into account the representation and generation of code edits. Based on code pre-trained models, code editing pre-trained models use the following two ways to consider code edits: MLM with code changes and DAE with controlled noises.
@@
CodeReviewer [7], CCT5 [11], and Coeditor [12] use MLM with code changes. The basic idea of these models is to use a sequence to represent a real-world code change. That is, the output of the diff tool when comparing the two code snippets is treated as the sequence for the code change. If there is a comment associated with the code change, the comment is also appended to the sequence. Note that the output of the diff represents the line-level differences on two code snippets. On top of the sequences representing code changes, CodeReviewer [7] and CCT5 [11] use the standard MLM for pre-training, where the masked parts can be tokens in the following types: tokens to be inserted, tokens to be deleted, tags generated by diff, and tokens in the comment. Different from CodeReviewer, CCT5 also adopts GPT-like pre-training tasks and adds data flow information into the sequence representing a code change. Similarly, Coeditor [12] first uses placeholders to mark the sections that need modification, then predicts the next diff information based on multiple prior line-level code changes. Our DivoT5 differs from CodeReviewer, CCT5, and Coeditor in two folds. First, DivoT5 uses token-level difference information for pre-training, while CodeReviewer and CCT5 are based on line-level differences. Second, the pre-training tasks in DivoT5 aim to predict code edits on top of the old code (i.e., ensuring evolutionary direction of code editing), thus well aligned with downstream code-editing tasks. However, the pre-training tasks in CodeReviewer, CCT5, and Coeditor do not correspond to any code-editing tasks. For example, the prediction of masked tokens when pre-training CodeReviewer and CCT5 may use information in both code snippets before and after change.
@@
CoditT5 [8] and CodeEditor [10] use DAE with controlled noises. The basic idea of these models is to control noise generation in order that the generated noises can represent the probabilities real-world changes. In particular, CoditT5 [8] utilize the CodeSearchNet [24] dataset to design a denoising task for pre-training. Specifically, they calculate the probabilities of inserting, deleting, and modifying tokens in real-world code changes, and use these probabilities to artificially perturb the code. Then the standard DAE tasks are adopted for pre-training on top of the perturbed code. Similar to CoditT5, CodeEditor [10] is also based on code perturbed with probabilities from real world. However, the perturbed code is not directly used in pre-training. CodeEditor utilizes an existing code-generation model [16] to further process the perturbed code, and the processed code is then used for pre-training. Thus, the pre-training may well focus on the inserted noises that can hardly be handled by existing models. Our DivoT5 differs from these models in that DivoT5 is based on predicting real-world code edits while CoditT5 and CodeEditor are based on removing artificially inserted noises. Note that, although CoditT5 and CodeEditor use real-world probabilities to control the insertion of noises, the perturbed code may still largely deviate from real-world human-written code. Furthermore, DivoT5’s pre-training tasks also incorporate noise to perturb the old code, but the aim is to help DivoT5 learn to predict code edits in the presence of noise.

### v2

@@ Abstract
Code pre-trained models have shown promising effectiveness in various software engineering tasks. Among these tasks, many tasks are related to software evolution and/or code editing. However, existing code pre-trained models often overlook the real-world code editing data and the evolutionary nature of the editing process. In this paper, to simulate the step-by-step code editing process of human developers, we propose DivoT5, a pre-trained model based on directional diffusion at the data level. In DivoT5, we adopt two categories of pre-training tasks. The first category is mask and denoising tasks augmented with a diffusion direction representing code evolution. That is, we first apply a noising process to the code snippets before evolution, and then ask the pre-training process to restore the snippets with noise into the code snippets after evolution. The second category is tasks aiming to reinforce the evolutionary direction. That is, we first generate various intermediate versions for each pair of snippets before and after evolution, and then ask the pre-training process to transform the intermediate versions into the snippet after evolution for each pair. We evaluate DivoT5 for two code-editing scenarios (including a number of tasks) and one non-editing scenario using four downstream tasks. For each downstream task, we fine-tune the pre-trained DivoT5 on multiple corresponding datasets and evaluate its effectiveness across diverse scenarios Our experimental results show that DivoT5 achieves state-of-the-art (SOTA) performance on most tasks in comparison to models of the same scale (220M), large-scale (770M, 6.7B) models in fine-tuning, and billion-scale (6.7B, 8B, ChatGPT) instruct models in few-shot settings. For one code-editing task (i.e., CodeReview in NL-based CodeRefinement task), DivoT5 pre-trained on top of CodeT5-small (60M) can even outperform CodeT5-base (220M) and other pre-trained models with 220M parameters except for DivoT5 pre-trained on top of CodeT5-base (220M).
@@ I Introduction
The code of software systems often requires continuous maintenance and updates to enhance functionality, adapt to new requirements, and fix vulnerabilities [lientz1978softwaremainten, bennett2000softwareevolution, tufano2021codereview, recoder2021, just2014defects4j], which are collectively termed “code evolution”. Throughout the code evolution process, many tasks require generating new code by editing existing old code, such as code refinement [tufano2019coderefine, codereviewer, tufano2021codereview, coditt52022, twinxsql] and bug fixing [tufano2019bug2fix, tufano2019learning_bugfixing]. These tasks related to code editing constitute an essential part of the code evolution process. To address tasks related to code editing, many studies have utilized pre-trained code models, such as CodeT5 [codet52021], given their proven ability to effectively encode the semantic information of code. However, directly applying these pre-trained models to code editing tasks often fails to achieve satisfactory performance because they do not adequately consider the contextual information (i.e., the relationship between the old and new versions of code during the evolution process). This context includes the semantic alignment across code versions and serves as critical guidance for accurately identifying what should be modified, retained, or extended.
@@
Recently, several studies [coditt52022, li2023codeeditor, codereviewer, lin2023cct5] propose pre-training tasks specifically designed for code editing tasks and learning contextual information of code evolution by fine-tuning pre-trained models. These models consider code editing through two ways: masked language modeling (MLM) with code changes [codereviewer, lin2023cct5] and denoising auto-encoding (DAE) with controlled noises [li2023codeeditor, coditt52022]. The first way represents code changes as the output sequence of the diff tool and applies the MLM pre-training tasks to understand the semantics of code changes. This way has shown promising results in understanding code changes. However, in real development scenarios, code evolution typically involves generating new code based on existing old code, rather than understanding the code change. Therefore, using code changes as input to generate corresponding descriptions (e.g., commit messages [lin2023cct5]) or classification labels (e.g., whether the code improves [codereviewer]) does not align well with such development scenarios for code generation. Moreover, the granularity they adopt is typically at the code line level [coeditor_iclr2024], as produced by the diff tool. However, capturing fine-grained changes at the token level is essential for accurately modeling developer intent and generating precise code modifications [edit_emperical, edit_emperical_2]. This also posing a challenge in considering fine-grained changes (i.e., token-level) during the generation process (i.e., the first challenge lies in reflecting fine-grained edits akin to those made by human developers). In the second way, noise is typically introduced to the old code based on the probability distribution derived from real code changes in downstream datasets, and the model employs the DAE-like pre-training tasks to restore the old code from the noisy input [coditt52022]. However, this noise-disrupted code only captures the general likelihood of edits, without accounting for the fact that different pieces of old code should evolve into contextually appropriate new versions. In other words, the editing process is not merely about reflecting the overall probability of code being edited, but about transforming code in ways that reflect the specific intent and semantics embedded in each unique context. Consequently, there exists a substantial gap that existing approaches often overlook between disrupted code and the imperfect, context-rich code written by real developers.  (i.e., the second challenge lies in leveraging the context of code edits made by human developers). In the development scenarios, code evolution typically involves a series of modifications to one file, making it a multi-step editing process rather than a one-off effort. Instead of focusing on only what the code changes are or what the actual change probabilities might be, how the code is gradually edited to achieve the final evolutionary goal should be also be considered. However, existing code editing pre-trained models only utilize the before-and-after code pairs directly, neglecting the gradual evolution of code in real-world scenarios and its characteristic of incremental editing  (i.e., the third challenge lies in capturing the incremental nature of code editing performed by human developers.).
@@
In this paper, we introduce DivoT5, a diffusion-style pre-trained model. To tackle the above challenges, the key idea of DivoT5 is to simulate the incremental nature of human code editing in the pre-training step. It draws inspiration from diffusion models, which excel at handling incremental transformations in continuous data domains, generating high-quality and diverse images or audio through a gradual multi-step denoising process [diffapp_ho2020denoising, difflm_ARDiff_wu2024ar, difflm_DiffLM_li2022diffusion]. Similarly, code editing involves making a series of iterative changes to fulfill the final requirements, making it naturally aligned with the step-by-step refinement principles of diffusion models. However, DivoT5 does not directly implement diffusion processes. It utilizes its data generation principles and focuses on simulating the gradual code editing steps that developers typically follow, allowing it to produce realistic and contextually relevant code modifications. To distinguish it from traditional diffusion methods, we refer to this approach as “directional diffusion.”
@@
Figure 1 illustrates the schematic of directional diffusion. We set the starting point (i.e., $X_{T}$) of the directional diffusion strategy as the old code, design three mask strategies as introducing artificial noise (i.e., $X_{T+1}$) to understand the evolutionary direction, and treat the intermediate states (i.e., $X_{t}$) of the editing process as evolutionary noise that reinforces the direction of evolution. Specifically, $X_{T+1}$ to $X_{T}$ represents the process of removing artificial noise (e.g., replacing $p_{1}$ and $v_{1}$ to $|unk|$ in Figure 1), while $X_{t}$ to $X_{t-1}$ represents the process of removing evolutionary noise (e.g., modifying the $f_{11}$ in the Figure 1). To fit the fine-grained editing scenarios, DivoT5’s pre-training takes the old code as input and incorporates token-level diff information when generating the complete new code. Rather than relying on line-level diff information to understand code changes, DivoT5 leverages token-level diff information to help the model discern which parts of the code need modification and which parts should remain unchanged (i.e., a mask strategy applied from $X_{T}$ to $X_{T+1}$ in Figure 1). To leverage and enrich the context of the real-world code edits, we design two additional mask strategies based on real code change histories. Learning from real code changes means that the model needs to automatically update old versions of code (which may contain bugs or require new features) to new versions that meet new requirements. These two masking strategies (applied from $X_{T}$ to $X_{T+1}$ in Figure 1) simulate potential issues with code quality or completeness in real code context by introducing varying degrees of random artificial noise into the old code. This requires the model to understand the context of the specific code and make modifications based on the provided guidance, rather than merely restoring the original version from the corrupted code. To learn the characteristics of incremental code evolution, we design a novel evolutionary direction reinforcement strategy by utilizing intermediate data in the code editing process. Intermediate versions (i.e., $X_{t}$ ($0<t\leq T$) in Figure 1) of the code that is not fully modified are seen as still containing evolutionary noise. We utilize intermediate data from the incremental evolution process to train the DivoT5 model, providing clear directional and more detailed guidance information to reinforce the evolutionary direction. Due to the limitations of applying existing diffusion models to discrete textual data, which typically use multi-step denoising at continued embedding levels and whose performance is constrained by the mapping between continuous embeddings and discrete texts, we apply the above strategies at the data level (i.e., conducting the diffusion process on data rather than embeddings). This model ensures the fluency and reliability of the generated results by directing each version towards the final version ($X_{0}$), even with the addition of artificial noise (i.e., constructing training samples $X_{t}$, $X_{0}>$ and generating $X_{0}$ from any given $X_{t}$, where $0<t\leq T+1$). In other words, we adopt the encoder-decoder training framework [codet52021] while incorporating the diffusion concept to emphasize the evolutionary direction of code editing. Rather than based on the full diffusion model structure, we draw on its style to enhance step-by-step refinement, ensuring the same computational efficiency as the encoder-decoder model [codet52021, coditt52022, codereviewer] during training and inference.
@@
To evaluate DivoT5, we fine-tune DivoT5 on three scenarios (i.e., NL-guided code editing scenario, code-only editing scenario, and non-editing scenario) using eight benchmark datasets. The evaluation results demonstrate that DivoT5 outperforms the existing state-of-the-art (SOTA) code pre-trained models for all eight datasets. For example, DivoT5-base (220M) achieves a significant improvement in exact match performance compared to the CodeT5-base (220M) model, exceeding a 28.87% increase on the CodeReview dataset (from 34.46% to 44.41%). Furthermore, DivoT5-base (220M) significantly outperforms the fine-tuned CodeT5-large (770M) model and billion-scale models like DeepSeekCoder-6.7B-instruct, Meta-Llama-3.1-8B-instruct, and ChatGPT-gpt-3.5-turbo in few-shot inference. The small version of DivoT5 (60M) also achieves a good performance in this task, surpassing CodeT5-base (220M) by 5 percentage points (from 34.46% to 40.16%) in the exact match metric. In addition, we evaluate the effectiveness of each pre-training task. The ablation results demonstrate the contributions of the designed pre-training tasks to the final performance.
@@ II-A Code-Editing Tasks and Code-Editing Related Tasks
During the daily development process, code editing is necessary to maintain the overall functionalities of a software project. In this paper, we differentiate two categories of tasks that are associated with code editing. The first category is tasks involving an editing process to change a code snippet, and we refer to this category as code-editing tasks [tufano2019coderefine, tufano2021codereview, li2023codeeditor, coditt52022, twinxsql]. The second category is tasks that do not directly edit code but may use code-editing information for other purpose, and we refer to this category as code-editing related tasks [lin2023cct5, codereviewer]. Obviously, it is helpful to automate both categories of tasks during the code maintenance process.
Based on whether natural language guidance is required during the code-editing process, we can further divide code-editing tasks into two types. We denote the first type as the NL-guided code editing scenario, such as code refinement [tufano2021codereview, coditt52022, codereviewer]. Developers provide suggestions for improvement by inspecting and evaluating the code, and the natural language represents the direction for code enhancement. The new code should align with this direction. We denote the second type as the code-only editing scenario, such as code refinement [codexglue2021]. For the code refinement task, the new code should address the defects present in the old code. In this paper, we primarily consider these two types of code-editing tasks.
@@ II-B Code Pre-trained Models
When applying existing pre-training techniques [bert2018, gpt2018, lewis2019bart, t52020, scis_deeplearning, grammarcoder] to code data, there are two common pre-training tasks, which are referred to as masked language modeling (MLM) [codebert2020, codet52021] and denoising auto-encoding (DAE) [plbart2021]. The basic MLM technique refers to randomly masking a certain proportion of tokens in the input and then predicting the masked tokens. CodeBERT [codebert2020] and GraphCodeBERT [graphcodebert2020] apply the MLM technique to its code data to enhance the model’s understanding of code. CodeT5 [codet52021] employs the MLM technique to sequence-to-sequence generation tasks (referred to as SMLM) to enhance the model’s ability to generate code. Additionally, to recognize identifiers within the code, CodeT5 also introduces a variant of SMLM called IMLM (Identifier Masked Language Modeling). In IMLM, only the identifiers in the source code are masked, while the rest of the tokens are left unchanged, allowing the model to generate code while specifically paying attention to the discernment of identifiers. DAE is a self-supervised learning approach that enhances model generalization by introducing noise to the input and training the model to reconstruct the original input. In PLBART [plbart2021], DAE is introduced during the pre-training stage to train the model in code understanding and generation capabilities. Specifically, PLBART incorporates three noise functions: token masking, token deletion, and token infilling. These functions are applied with a certain probability to disrupt the old code.
@@ II-C Code Editing Pre-trained Models
Code editing pre-trained models are typically initialized from existing code pre-trained models but are further adapted to specialize in code editing tasks. These models aim to better capture the structure and semantics of code transformations by introducing edit-aware pre-training objectives. Based on code pre-trained models, code editing pre-trained models use the following two ways to consider code edits: MLM with code changes and DAE with controlled noises.
@@
CodeReviewer [codereviewer], CCT5 [lin2023cct5], and Coeditor [coeditor_iclr2024] use MLM with code changes. The basic idea of these models is to use a sequence to represent a real-world code change. That is, the output of the diff tool when comparing the two code snippets is treated as the sequence for the code change. If there is a comment associated with the code change, the comment is also appended to the sequence. Note that the output of the diff represents the line-level differences on two code snippets. On top of the sequences representing code changes, CodeReviewer [codereviewer] and CCT5 [lin2023cct5] use the standard MLM for pre-training, where the masked parts can be tokens in the following types: tokens to be inserted, tokens to be deleted, tags generated by diff, and tokens in the comment. Different from CodeReviewer, CCT5 also adopts GPT-like pre-training tasks and adds data flow information into the sequence representing a code change. Similarly, Coeditor [coeditor_iclr2024] first uses placeholders to mark the sections that need modification, then predicts the next diff information based on multiple prior line-level code changes. Our DivoT5 differs from CodeReviewer, CCT5, and Coeditor in two folds. First, DivoT5 uses token-level difference information for pre-training, while CodeReviewer and CCT5 are based on line-level differences. Second, the pre-training tasks in DivoT5 aim to predict code edits on top of the old code (i.e., ensuring evolutionary direction of code editing), thus well aligned with downstream code-editing tasks. However, the pre-training tasks in CodeReviewer, CCT5, and Coeditor do not correspond to any code-editing tasks. For example, the prediction of masked tokens when pre-training CodeReviewer and CCT5 may use information in both code snippets before and after change.
@@
CoditT5 [coditt52022] and CodeEditor [li2023codeeditor] use DAE with controlled noises. The basic idea of these models is to control noise generation in order that the generated noises can represent the probabilities real-world changes. In particular, CoditT5 [coditt52022] utilize the CodeSearchNet [csn2019] dataset to design a denoising task for pre-training. Specifically, they calculate the probabilities of inserting, deleting, and modifying tokens in real-world code changes, and use these probabilities to artificially perturb the code. Then the standard DAE tasks are adopted for pre-training on top of the perturbed code. Similar to CoditT5, CodeEditor [li2023codeeditor] is also based on code perturbed with probabilities from real world. However, the perturbed code is not directly used in pre-training. CodeEditor utilizes an existing code-generation model [codexglue2021] to further process the perturbed code, and the processed code is then used for pre-training. Thus, the pre-training may well focus on the inserted noises that can hardly be handled by existing models. Our DivoT5 differs from these models in that DivoT5 is based on predicting real-world code edits while CoditT5 and CodeEditor are based on removing artificially inserted noises. Note that, although CoditT5 and CodeEditor use real-world probabilities to control the insertion of noises, the perturbed code may still largely deviate from real-world human-written code. Furthermore, DivoT5’s pre-training tasks also incorporate noise to perturb the old code, but the aim is to help DivoT5 learn to predict code edits in the presence of noise.